Using the z table (table E) find the critical value (or values) for the case where a=0.01, two tailed test.

Using the z table (table E) find the critical value (or values) for the case where a=0.05, right tailed test

Using the z table (table E) find the critical value (or values) for the case where a=0.02, left tailed test.
Answer #1

Solution :

1)

\alpha = 0.01

\alpha / 2 = 0.01 / 2 = 0.005

Z\alpha/2 = Z 0.005 = 2.576

Critical value = 2.576

2)

Critical value = 1.645

3)

Critical value = -2.054
Answer #2

α = 0,05